OpenStudy (anonymous):

Keisha's parents want to save $20,000 in her college savings account over the next 15 years. They have $8,000 to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth.

OpenStudy (amistre64):

simple interest is non compunding so it works on th einitial principle alone; the need 12000 more dollars 12000 = 8000 r 15 r = 1/10 = .1 or 10%

OpenStudy (anonymous):

20000 = 8000 * (1 + x) ^15 x = 6.3%